OpenStudy (anonymous):

Differentiate and Simplify: f(x)= e^(2x) * (x-1)

OpenStudy (anonymous):

\[f(x)= e^{2x}(x-1)\]

OpenStudy (anonymous):

Use the product rule. \[f'(x)=e^{2x}*1+(x-1)*2e^{2x}\] \[f'(x)=e^{2x}*2e^{2x}(x-1)\] \[f'(x)=e^{2x}(2x-2)\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

f′(x)=e2x∗1+(x−1)∗2e2x f'(x)= e2x(1+2x-2) f'(x)=e2x(2x-1)